Last year, the mean waiting time for the number 14 bus was 11.4 minutes.  This year, a random sample of 12 waits for the bus yielded the following waiting times, in minutes.  Use the Wilcoxon signed-rank test to perform the required hypothesis test. Last year, the mean waiting time for the number 14
Answer #1
Values ( X ) \Sigma (X_{i} - \bar{X})^{2}
10.2 12.3082
14.6 0.7951
6 59.4179
20.5 46.1272
2.7 121.1827
23.6 97.8457
14.1 0.1534
8.6 26.0947
30.6 285.3295
4.7 81.1495
12 2.9183
16.9 10.1869
Total 164.5 743.5091

Mean \bar{X} = \Sigma X_{i} / n
Y = 164.5/12= 13.7083
Standard deviation S_{X} = \sqrt{\Sigma (X_{i} - \bar{X})^{2}/n-1}
sx = V743.509 1/12-1-8.2214

To Test -

H0 :-  μ = 11.4

H1 :-  μ>11.4

α 0.01

Test Statistic :-
t = ( \bar{X} - \mu ) / (S /\sqrt{n})
t = (13.7083-11.4)/(8.2214/v/12)
t = 0.9726

Test Criteria :-
Reject null hypothesis if t > ta,n-1

Critical value  ta,n-l t0.01 12-1 2.718

ta,n-1 = 0.9726 < 2.718
Result :- Fail to reject null hypothesis


Decision based on P value
P - value = P ( t > 0.9726 ) = 0.1758
Reject null hypothesis if P value < α 0.01 level of significance
P - value = 0.1758 > 0.01 ,hence we fail to reject null hypothesis
Conclusion :- Fail to reject null hypothesis

At 1% significance level, the data does not provide sufficient evidence to support the claim that mean waiting time this year is higher.